1. 2-way Set Associative Cache Memory
Consider a hypothetical machine with 1K words of cache memory. They are in

two-way set associative organization, with cache block size of 128 words, using LRU replacement algorithm. Suppose the cache hit time is 9ns, the time to transfer the first word from main memory to cache is 50ns, while subsequent words require 10ns/word.

Consider the following read pattern (in blocks of 128 words, and block id starts from 0):

    1 2 3 5 6 2 3 4 9 10 11 6 3 6 1 7 8 4 5 9 11 1 2 4 5 12 13 14 15
    13 14

and assume each block contains an average of 48 references.

  1. (a) What is the cache miss penalty (i.e., time to transfer one block of data from main memory to cache memory)?
Answer #1

a) Cache miss penalty:

Number of words in a cache block size = 128 words

Time to transfer the first word from main memory to cache in case of a miss = 50 ns

Time to transfer rest 127 words = 10*127 = 1270 ns

Therefore, time to transfer one block of data from main memory to cache memory( cache miss penalty ) = 1270 + 50 = 1320 ns.